Circular No. 6367

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) J. Lecacheux, Meudon Observatory; E. Frappa, Planetarium of Saint-

Etienne; P. Laques, Pic du Midi Observatory; F. Colas, Bureau des Longitudes, Paris; and L. Jorda, Observatoire de Meudon and Laboratoire d’Astronomie Spatiale, Marseille, report: “Monitoring has continued at the Pic-du-Midi 1.05-m telescope of fragments moving away from the comet’s nucleus. Small fragments are difficult to identify from night to night, but measurement of their velocity is possible in a few hours. Within 10 000 km of the nucleus, their velocity dr/dt along the extended radius vector is typically on the order of 5R m/s, where R is expressed in units of 1000 km. The major fragment observed by Tozzi et al. on Mar. 25.93 UT (IAUC 6357), and later by the Hubble Space Telescope on Mar. 26.06 (IAUC 6363), had a velocity of 11.5 m/s at that time, according to our measurements; we located it subsequently on Mar. 28.9 at R = 4800 km (diffuse appearance), then at R = 7450 km on Mar. 30.9 (stellar appearance). Its magnitude brightened from 16 to 14 between Mar. 30.79 and 30.86. It seemed diffuse again and was of mag 13 at R = 9150 km on Mar. 31.81.”

W. M. Harris, F. Scherb, J. W. Percival, and K. Nordsieck, University of Wisconsin; K. Honeycutt, Indiana University; and B. E. A. Mueller, D. Harmer, and A. Macdonald, National Optical Astronomy Observatories, report on multiple observations of a symmetric conical structure extending behind comet C/1996 B2 in images taken with the 3.5-m WIYN telescope at Kitt Peak: “The conical structure has been consistently observable in images taken since Mar. 24 using two filters that isolate bright, resonant CN and OH neutral-gas emissions. It has not been detected in WIYN images targeting CO^{+} and H_{2}O^{+} ion emission lines, at narrowband continuum points at 683.2 and 487.5 nm, or in the broadband R. The shape of the feature is most reminiscent of a 3-dimensional analog to a ship’s wake. It emanates from the near-nuclear vicinity with an 80^{o}–120^{o} opening angle and is visible as diffuse, low-contrast emission out to a distance of roughly 2000 km. At this point, a sharp increase in brightness is observed as a partial arc surrounding the primary dust lane behind the nucleus. This may represent a line-of-sight enhancement of the optical depth of gas in the feature, or possibly a localized gas-density increase. Emission is also detectable from beyond the arc, but the contrast is again very low. An image showing the conical structure is available at URL

http://www.noao.edu/noao/comets/hy/wiynhy.html.”

1996 April 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green